# Inverse design of potential metastructures inspired from Indian medieval architectural elements

### v2

@@ Abstract
In this study, we immerse in the intricate world of patterns, examining the structural details of Indian medieval architecture for the discovery of motifs with great application potential from the mechanical metastructure perspective. The motifs that specifically engrossed us are derived from the tomb of I’timad-ud-Daula, situated in the city of Agra, close to the Taj Mahal. In an exploratory study, we designed nine interlaced metastructures inspired from the tomb’s motifs. We fabricated the metastructures using additive manufacturing and studied their vibration characteristics experimentally and numerically. We also investigated bandgap modulation with metallic inserts in honeycomb interlaced metastructures. The comprehensive study of these metastructure panels reveals their high performance in controlling elastic wave propagation and generating suitable frequency bandgaps, hence having potential applications as waveguides for noise and vibration control. Finally, we developed novel AI-based model trained on numerical datasets for inverse design of metastructures with a desired bandgap.
@@
Keywords: Metamaterials, Metastructures, Interlaced metastructures, Hierarchical Metastructures, Fractals, Honeycomb, Elastic wave control, Artificial Intelligence, Multihead geometric attention, Multiscale residual network, Inverse design.
Introduction
@@
Architectural intricate patterns have a long history, with certain examples dating back to the 4th century [1]. However, these early forms were the outcome of creative heuristic advancements of pattern generation in human civilization. It wasn’t until the late 19th century, when fundamental physics became better understood, that the development of systematic engineering principles made it possible to explore counter-intuitive physical characteristics of some patterns in periodic structures, generally known as metamaterials and metastructures. Today, it is often proposed as a panacea for controlling attenuation and wavefront divergence related to elastic and electromagnetic wave propagation and there exists a vast literature of which are based on meticulously designed lattice structures with special geometric patterns [2, 3, 4, 5, 6]. These engineered lattices may introduce exotic mechanical and electromagnetic properties, manipulating the behavior of electromagnetic waves [7], elastic waves as well as sound waves [8, 9, 10, 11]. By carefully designing these lattices, scientists can achieve extraordinary effects like negative stiffness [12], negative mass [13], and negative Poisson’s ratio [14]. The study of lattice geometries in metamaterials opens a pathway to myriad innovative applications ranging from optics to acoustics, wearable devices [15, 16], advanced space structures [17, 18], next-gen automobiles [19, 20], robotic manipulators [21], bionic grippers, [22] and anti-seismic structures [23, 24]. These metamaterials exhibit characteristics such as high specific stiffness [25, 26, 27, 28], negative Poisson’s ratio [29, 30], negative thermal expansion coefficient [31], negative reflection [32, 33], wave alteration [34], sub-wavelength bandgap [35, 36] and distinctive electro-dynamic phenomena [37, 38]. The generation of a large range of novel periodic structures at different length scales has ushered in phenomenal developments of metamaterials.
Figure 1: The Tomb of Itmad-ud-Daula is a Mughal mausoleum in the city of Agra in the Indian state of Uttar Pradesh (27.1929 N, 78.0310 E). The architecture of this monument is unique in terms of crafting which uses a combination of pietra-dura and calligraphy-based decorative art techniques. It is considered to be the pioneer of the creation of funerary tombs in Mughal-India which was subsequently magnified in the Taj Mahal. However, with its marble latticework and exquisite carvings, the tomb is considered as a work of art even more delicate than the Taj Mahal. (b to d). Consequently, the latticework is depicted through the carved panels of calligraphic design. The walls, floors, and niches are decorated with beautiful geometrical patterns. (e and f). Here, we have specifically focused on the Jali screens at the centre of the niches all around the tomb. They were used extensively during the Mughal reign for elegance and their ability to facilitate the entry of natural light and air inflow. Often it is also used as part of a passive cooling system. (g) shows the corresponding structural models (LS1-LS9).
@@
Mechanical properties of simple metastructures such as honeycombs with positive and negative re-entrant angles have been extensively studied through experimental avenues, numerical simulations, and mathematical analyses [39, 40, 41, 42, 43, 44]. In recent years, there has been a growing interest in developing engineered materials using a range of 3D and 4D additive manufacturing techniques [45, 46]. Recent advancements in precise additive manufacturing techniques at the microscale have further enabled the fabrication of new lattices with complex topology and interlaced geometry, which are challenging to achieve using traditional subtractive manufacturing methods. Such metastructures have distinct behaviors when subjected to external stimulus such as mechanical excitation [47, 48, 49, 50, 51, 52, 53].
Geometric design and pattern have played a significant role in the spiritual traditions of most of the religions. For instance, concentric circles and triangles are widely used as sacred patterns in the Hindu and Buddhist ‘Mandala’; while rectangles and squares are preferred in the ‘Feng Shui’ design of Chinese spiritual tradition. The construction of medieval European cathedrals was often based on the ‘Latin Cross’. Fractal geometries were the most popular design used in the construction of various Hindu temples. The pinnacle of embedding geometric motifs is found in ‘Islamic Geometric Patterns’ which involve more complicated shapes such as hexagons, octagons, decagons, and multi-point stars [54]. One of the interesting applications of such motifs in tessellated form is called ‘Mashrabiya’ which is known to be originated in Baghdad during the 12th-century Abbasid period. Its interlaced arrangements, which emerged from simplicity and unity of natural as well as geometric motifs, exhibit some fundamental relationships that become in-turn hierarchical or fractal in complex architectures. The architectural applications of ‘Mashrabiya’ had rapidly developed in the medieval Persian and Indian civilizations due to the influence of Islamic religion and mysticism. Such patterns, popularly known as ‘Jali’ in India, were crafted not only for aesthetic purposes but also for their utility in terms of transmitting indirect sunlight, capturing fresh air, passive cooling, maintaining communication with privacy, etc. (some of the patterns are shown in Fig. 1(b-f)). The first instance of developing such patterns in India was found in the Alai Darwaza in the Qutb Minar complex of Old Delhi. The tomb of Itimaduddaula, a magnificent testament to the Mughal style enriches Agra’s vibrant tapestry. Considered as the inspiration for the Taj Mahal, the tomb also known as the ‘Baby Taj Mahal’, is a stunning example of Mughal architecture (Fig. 1 (a)). This mausoleum displays a wonderful variety of geometric arabesques. The strong influence of the Persian style coupled with numerous Indian motifs is visible in this structure. The related lattice structures can be classified based on different unit cell types, considering angle, arrangement, and symmetry (Figure 2(a)). We investigate various interlacing arrangements of these unit cells (Figure 2(b)) and demonstrate how different arrangements can be used to generate complete metastructures (Figure 2(c)). These metastructures provide a foundation for further studies evaluating their performance in controlling elastic wave propagation and vibration. Further, we have developed an AI-based forward analysis model to predict the transmissibility of a metastructure. Also, an AI-based inverse design of metastructure is developed to design new lattices for the desired bandgap creating end-less possibilities of generating suitable patterns focused to an application [55, 56, 57, 58, 59, 60].
@@
Figure 3: Schematic diagram illustrates latticed mechanical metastructures featuring different interlaced lattice structures (LS1-LS9). Dynamic analysis of the metastructures is performed and mode shapes of lattices at a frequency corresponding to minimum force transmissibility for in-plane excitation are shown. The attenuations corresponding to each geometry are obtained by using the transmissibility plot for each structure. We have highlighted distinct attenuation zones within the transmissibility response of metastructures through green-shaded patches.
@@
Figure 4: Conducting a comparative analysis, this study investigates the impact of scaling (x, 1.5x, 2x) of the unit cell on the bandgap, with the resulting BG ratio illustrated in a bar diagram. The basic building blocks of unit and sub-unit cells are represented in the middle column of the figure. Subunit cells are used as an interlaced component for the primary structure from which each configuration is constructed. The patterns of latticed mechanical metastructures are 3D Printed using PLA having different primary lattice structures namely (a) Honeycomb-honeycomb interlaced (LS1) (b) Honeycomb interlaced with honeycomb pattern (LS2) (c) Honeycomb interlaced with a star (LS3) (d) Chiral interlaced structure (LS4) (e) Ring interlaced structure (LS5) (f) Fishtail structure (LS6) (g) Star-honeycomb structure (LS7) (h) Honeycomb-auxetic structure (LS8) (i) Honeycomb-auxetic interlaced structure (LS9).
@@
Results
Dynamic characterization and determination of the Bandgap
Selected latticed metastructures (represented in Fig. 3), generated based on architectural elements (described in Fig. 2) have an overall size of 204 mm x 58 mm and a thickness of 5 mm. The size was carefully chosen keeping in mind the manufacturing constraints as well as the requirement of minimum number of periodic units (greater than 7 units) in the tessellated structure to express the metamaterial behavior [13]. On both sides, 15 mm solid frames are appended with the latticed structure in the longitudinal direction (see Supplementary Fig. S1) for ease of clamping and excitation during dynamic testing. For each latticed structure, there is a frequency band in which the structure may work as an attenuator signifying the absence of wave propagation, and the same is highlighted in Fig. 3. In the middle column (Fig. 3), we have shown the related mode shapes obtained using Ansys corresponding to minimum force transmissibility within the attenuation zone. Different structures have different damping ratios depending on the lattice geometry. The damping ratio is calculated by the half-power bandwidth method (highest for LS1 and the lowest for LS9), which is shown in the supplementary Fig. S11.
Latticed metastructures with different sub-lattice structures are shown in Fig. 4. Figure 4(a-c) illustrates the primary unit cells as honeycomb sub-lattice, single honeycomb, circular pattern of honeycomb cells and star structure respectively. In Figure 4(d), the primary unit cell has a chiral structure with sub-lattices exhibiting a spider web-like pattern. Furthermore, Fig. 4(e) features a primary unit cell as a ring structure interlaced with other ring structures. Figure 4(f) is a fishtail structure. Fig. 4(g) shows the primary unit cell as a series of honeycombs arranged to form a star at the center of the structure. Figure 4(h, i) are combinations of honeycomb and auxetic structure. In the former, these structures are present in a series combination, while in the latter, the auxetic structure is interlaced with the honeycomb structure. The dimensions of the unit cells are detailed in Supplementary Table S1. After creating all the metastructures of specified dimensions, we scaled them up by 1.5 and 2 respectively to capture the scale effect of unit cells in the wave transmission pattern. The scaling effect is illustrated with the help of the Band Gap (BG) ratio in the bar diagram (Fig. 4).
Performance measurement of lattice with resonators

### v3

# AI-driven Inverse Design of Band-Tunable Mechanical Metastructures for Tailored Vibration Mitigation

@@ Abstract
On-demand vibration mitigation in a mechanical system needs the suitable design of multiscale metastructures, involving complex unit cells. In this study, immersing in the world of patterns and examining the structural details of some interesting motifs are extracted from the mechanical metastructure perspective. Nine interlaced metastructures are fabricated using additive manufacturing, and corresponding vibration characteristics are studied experimentally and numerically. Further, the band-gap modulation with metallic inserts in the honeycomb interlaced metastructures is also studied. AI-driven inverse design of such complex metastructures with a desired vibration mitigation profile can pave the way for addressing engineering challenges in high-precision manufacturing. The current inverse design methodologies are limited to designing simple periodic structures based on limited variants of unit cells. Therefore, a novel forward analysis model with multi-head FEM-inspired spatial attention (FSA) is proposed to learn the complex geometry of the metastructures and predict corresponding transmissibility. Subsequently, a multiscale Gaussian self-attention (MGSA) based inverse design model with Gaussian function for 1D spectrum position encoding is developed to produce a suitable metastructure for the desired vibration transmittance. The proposed AI framework demonstrated outstanding performance corresponding to the expected locally resonant bandgaps in a targeted frequency range.
keywords:
@@
Metamaterials , Interlaced metastructures , Fractals , Elastic wave control , Geometric feature extraction , Multihead FEM-inspired spatial attention , Multiscale residual network.
journal: Mechanical Systems and Signal Processing
[…]
organization=Department of Mechanical Engineering, Indian Institute of Technology Kanpur,postcode=208016, country=India
[…]
organization=Department of Precision and Microsystems Engineering, Delft University of Technology,postcode=2628 CD, country=Netherlands
[…]
organization=Department of Mechanical Engineering, Motilal Nehru National Institute of Technology Allahabad,city=Prayagraj, postcode=211004, country=India
1 Introduction
@@
Metamaterials are emerging from varied research domains like electromagnetic, mechanical, and acoustic applications having unique tailored properties that are primarily dependent on their structural geometry rather than intrinsic material properties. It wasn’t until the late 19th century, when fundamental physics became better understood, that the development of systematic engineering principles made it possible to explore counter-intuitive physical characteristics of some patterns in periodic structures, generally known as metamaterials and metastructures. Today, it is often proposed as a panacea for controlling attenuation and wavefront divergence related to elastic and electromagnetic wave propagation and there exists a vast literature based on meticulously designed lattice structures with special geometric patterns [1, 2, 3, 4, 5, 6]. Generally, these structures have sub-wavelength features designed in the form of lattices on a smaller scale than the wavelength of external physical stimuli they interact with. These engineered lattices may introduce exotic mechanical and electromagnetic properties, manipulating the behavior of electromagnetic waves [7], elastic waves as well as sound waves [8, 9, 10]. As metastructures have structural-dependent properties, hence by carefully designing these lattices, scientists can achieve customizable responses with extraordinary effects like negative stiffness [11, 12, 13], negative mass [14], negative Poisson’s ratio [15, 16], negative thermal expansion coefficient [17], negative reflection [18, 19] for wave alteration [20], sub-wavelength bandgap [21, 22, 23] and distinctive electro-dynamic phenomena [24, 25, 26]. The study of lattice geometries in metamaterials opens a pathway to myriad innovative applications ranging from optics to acoustics such as wearable devices [27, 28], advanced space structures [29, 30], next-gen automobiles [31, 32], high precision robotic manipulators [33], bionic grippers, [34], anti-seismic structures [35, 36], and vibration sensing [37]. Mechanical properties of metastructures designed with simple unit cells such as honeycombs with positive and negative re-entrant angles, chiral, and kagome have been extensively studied through experimental avenues, numerical simulations, and mathematical analyses [38, 39, 40, 41, 42, 43, 44]. The generation of a large range of novel periodic structures at different length scales has ushered in phenomenal developments of metamaterials as depicted in Fig. 1.
@@
Architectural patterns in metamaterials can also play a pivotal role in manipulating electromagnetic and acoustic waves. Intricate architectural patterns have a long history, with certain examples dating back to the 4th century [45]. These early forms were the outcome of creative heuristic advancements of pattern generation often driven by culture and religion. Geometric design and pattern have played a significant role in the spiritual traditions of many religions. Considered as the inspiration for the Taj Mahal, the tomb of Itmad-ud-Daula, also known as the ‘Baby Taj Mahal,’ is a stunning example of Indian medieval architecture as shown in Fig. 2 (a). This mausoleum displays a wonderful variety of geometric arabesques. Such intricate patterns, popularly known as ‘Jali’ in India, were crafted for aesthetic purposes and also as a functional component in passive cooling and ventilation. Some of the patterns are shown in Fig. 2(b-f) with the corresponding metastructural models shown in Fig. 2(g). The related lattice structures can be classified according to different types of unit cells, considering angle, arrangement, and symmetry (Fig. 1(a)). We investigate various complex/hybrid arrangements of these unit cells (Fig. 1(b)) and then demonstrate how different arrangements can be used to generate complete metastructures (Fig. 1(c)). These metastructures provide a foundation for further studies evaluating their performance in controlling elastic wave propagation and vibration.
When we delve into the realm of complex hybrid unit cell-based metastructure it requires advanced manufacturing techniques like 3D printing. Recent advancements in precise additive manufacturing techniques at the microscale have further enabled the fabrication of new lattices with complex topology and interlaced geometry, which are challenging to achieve using traditional subtractive manufacturing methods. Such metastructures have distinct behaviors when subjected to external stimulus such as mechanical excitation [46, 47, 48, 49, 50]. Additive manufacturing has the potential to fabricate precisely complex cellular and architectural metamaterials by manipulating multimaterials with controllable properties. The design and fabrication of smart material-based programmable mechanical metamaterials with 4D printing technology can change their properties dynamically in response to external stimuli. In recent years, there has been a growing interest in developing engineered materials using a range of 3D and 4D additive manufacturing techniques [51, 52]. Such metastructures with complex multiscale geometry are difficult to design according to a desired frequency response with a targeted bandgap. It may require numerous attempts of FEM-based simulations and design changes to achieve the targeted frequency response and, hence, may not be feasible in practice due to high computational costs. Therefore, efficient AI-based models are necessary that can predict the effective design of a metastructure with complex geometry and the desired frequency response.
In recent years, various attempts have been made to develop AI-based inverse design models that can generate complex multiscale mechanical metamaterials having tailored mechanical response and properties [53, 54, 55]. Deng et. al. [56] suggested a neural accelerated evolutionary strategy that can obtain a geometry of metamaterials with hinged quadrilaterals for a wide range of target nonlinear mechanical responses.
Figure 2: (a)The Tomb of Itmad-ud-Daula is a mausoleum in the city of Agra in the Indian state of Uttar Pradesh (27.1929 N, 78.0310 E). The architecture of this monument is unique in terms of crafting which uses a combination of pietra-dura and calligraphy-based decorative art techniques. (b to d). Consequently, the latticework is depicted through the carved panels of calligraphic design. The walls, floors, and niches are decorated with beautiful geometrical patterns. (e and f). Here, we have specifically focused on the Jali screens at the centre of the niches all around the tomb. They were used extensively for elegance and their ability to facilitate the entry of natural light and air inflow. Often it is also used as part of a passive cooling system. (g) Shows the corresponding metastructure models (LS1-LS9) inspired from these architectural elements.
In another similar work, Ha et. al. [57] developed a rapid inverse design method using generative machine learning to generate metamaterials having uniaxial compressive stress-strain response. Recently, Bastek et. al. [58] proposed video-denoising diffusion models that can produce materials with tailored complex target performance. Their AI-based model, having spatial and temporal attention, enables the inverse design of metamaterial with a highly non-linear stress-strain response that closely matches with the finite element simulations. These research works were dedicated to the inverse design of metamaterials with tailored nonlinear stress-strain properties. Similarly, various researchers have also investigated the AI-based inverse design of metamaterials and metasurfaces like non-periodic 3D architectures having predictable anisotropic properties[59], low-index-contrast structures on a silicon photonics platform [60] and antenna surface design [61]. All these studies have generally focused on the inverse design of unit cells of metamaterials or periodic surfaces for a tailored mechanical or electrical response.
The inverse design of mechanical metastructures having complex interlaced geometry for tailored frequency response in a defined attenuation bandwidth still remains a challenge due to hierarchical multiscale geometry. Therefore, through this study, we propose a FEM-inspired geometric attention mechanism that can learn the geometry having multiple scales in a metastructure and produce the corresponding frequency response. Further, with the help of the forward model, an inverse design model is developed that can generate a suitable metastructure for the desired vibration response. As mechanical metamaterials have structure-dependent properties hence, their inverse design can revolutionize various application domains with customizable responses by tuning mechanical properties such as stiffness deformation, flexibility, and vibration response. Hence AI-based inverse design of mechanical metamaterials can enable advanced functionalities in prosthetics, exoskeletons, and bone implant designs, dynamic and adaptive structure design in robotics, impact and shock energy absorption lightweight metastructures, and soft robotics [62].
2 Numerical Simulations
Finite element analysis-based simulation has been carried out using Ansys, employing the Ansys Parametric Design Language (APDL) solver for all 720 structures, and the host material is chosen as PLA (detailed properties given in Supplementary Table S3). The entire structure was discretized by using tetrahedron elements with a 2 mm element size, to ensure convergence for interlaced geometries. The smaller edges of the base structure were fixed, while the longer edges remained free. A dynamic force of amplitude equivalent to 10 N was applied to the base, and the response was calculated on the opposite phase. Modal analysis was performed to identify all modes up to 10 KHz frequency, followed by harmonic analysis to obtain the transmissibility. The mode superposition method was used to link the modal analysis results with the harmonic analysis input. Frequency response functions (FRFs) were generated over the frequency range of 0 to 10 KHz with a 10 Hz frequency resolution; this step size is preferable to accurately capture the response. A structural damping ratio of 0.001 was incorporated to account for the viscoelastic effect of PLA substrate.
2.1 Dynamic characterization and determination of the Bandgap
@@
Figure 3: Schematic diagram illustrates latticed mechanical metastructures featuring different interlaced lattice structures (LS1-LS9). Dynamic analysis of the metastructures is performed and mode shapes of lattices at a frequency corresponding to minimum force transmissibility for in-plane excitation are shown. In mode shapes, red indicates maximum deformation and blue indicates minimum deformation. The attenuation corresponding to each geometry is obtained by using the transmissibility plot for each structure. We have highlighted distinct attenuation zones within the transmissibility response of metastructures through green-shaded patches.
Selected latticed metastructures as shown in Fig. 3, generated based on architectural elements (described in Fig. 2) have an overall size of 204 mm x 58 mm and a thickness of 5 mm.
@@
Figure 4: Conducting a comparative analysis, this study investigates the impact of scaling (x, 1.5x, 2x) of the unit cell on the bandgap, with the resulting BG ratio illustrated in a bar diagram. The basic building blocks of unit and sub-unit cells are represented in the middle column of the figure. Subunit cells are used as an interlaced component for the primary structure from which each configuration is constructed. The patterns of latticed mechanical metastructures are 3D Printed using PLA having different primary lattice structures namely (a) Honeycomb-honeycomb interlaced (LS1) (b) Honeycomb interlaced with a honeycomb pattern (LS2) (c) Honeycomb interlaced with a star (LS3) (d) Chiral interlaced structure (LS4) (e) Ring interlaced structure (LS5) (f) Fishtail structure (LS6) (g) Star-honeycomb structure (LS7) (h) Honeycomb-auxetic structure (LS8) (i) Honeycomb-auxetic interlaced structure (LS9).
The size was carefully chosen keeping in mind the manufacturing constraints as well as the requirement of minimum number of periodic units (greater than 7 units) in the tessellated structure to express the metamaterial behavior [14]. On both sides, 15 mm solid frames are appended with the latticed structure in the longitudinal direction (see Supplementary Fig. S1) for ease of clamping and excitation during dynamic testing. For each latticed structure, there is a frequency band in which the structure may work as an attenuator signifying the absence of wave propagation, and the same is highlighted in Fig. 3. We have shown the related mode shapes obtained using Ansys corresponding to minimum force transmissibility within the attenuation zone. Different structures have different damping ratios depending on the lattice geometry. The damping ratio provided is calculated by the half-power bandwidth method (highest for LS1 and the lowest for LS9), which is in the supplementary Fig. S10.
Latticed metastructures with different sub-lattice structures are shown in Fig. 4. Figure 4(a-c) illustrates the primary unit cells as honeycomb sub-lattice, single honeycomb, circular pattern of honeycomb cells and star structure respectively. In Figure 4(d), the primary unit cell has a chiral structure with sub-lattices exhibiting a spider web-like pattern. Furthermore, Fig. 4(e) features a primary unit cell as a ring structure interlaced with other ring structures. Figure 4(f) is a fishtail structure. Fig. 4(g) shows the primary unit cell as a series of honeycombs arranged to form a star at the center of the structure. Figure 4(h, i) are combinations of honeycomb and auxetic structure. In the former, these structures are present in a series combination, while in the latter, the auxetic structure is interlaced with the honeycomb structure. The dimensions of the unit cells are detailed in Supplementary Table S1. After creating all the metastructures of specified dimensions, we scaled them up by 1.5 and 2 respectively to capture the scale effect of unit cells in the wave transmission pattern. The scaling effect as shown in Fig. 4 is illustrated with the help of the Band Gap (BG) ratio in the bar diagram.
2.2 Bandgap characteristics of honeycomb-based interlaced structures
Figure 5: (a) A comparison of bandgap at different frequency levels of different structures interlaced in the honeycomb structure. (b), (d) The comparison of static performance is obtained through analytical computations of different structures. The displacement control mechanism is set as 5 mm compression of the initial height of samples (c) Comparison of bandgap in different order (n=0, 1, 2) of the fractal structure.
Comparative analysis of the effect of different interlaced structures on bandgap shift (Fig. 5(a)) is carried out in various frequency ranges. The band gap occurs at a frequency band comparatively higher than that of the simple honeycomb structure. Notably, for the LS3 structure bifurcation in bandgap occurs, and for the last structure (LS9) multiple bandgaps occur both for the higher and lower frequency ranges. Furthermore, we have compared the bandgap range of various interlaced honeycomb unit cells (Fig. 5(b)) along with the force-deflection relationships. The force-deflection relationships are linear for all the patterns except for LS9.
Figure 5(c) illustrates the fractal structure of different order (n=0, 1, 2) and their corresponding bandgap shifts. In different structures, as the fractal order increases, the bandgap shifts to higher frequency ranges, and for further increase in fractal order multiple bandgaps occur. Furthermore, on the right side we have represented the force-deflection relationship for the fractal structure (see Fig. 5(d)). With the increase in the fractal order, the structure becomes stiffer, with 196.32% increase from zeroth to first-order fractal and 33.61% increase from first to second-order fractal. From this study, we can conclude that with increasing the structural complexity inside the structure, the bandgap starts to shift towards the higher frequency range, and for higher-order fractals, it also bifurcates which is verified by taking different fractal structures inside the honeycomb and explained in the Supplementary Fig. S8.
3 Experimental testing
Figure 6: (a) The experimental setup, accompanied by the Laser Doppler Vibrometer (LDV), is depicted in the schematic diagram. (b) 3D laser head with a controller is featured. (c) Power amplifier (d) Electro-dynamic shaker which is employed in setup (e)-(f) Excitation is provided from the base through an electrodynamic shaker, while the response is obtained by Laser Doppler Vibrometer focused on the top surface of the work-piece.
@@ 3 Experimental testing
The study utilizes a non-contact technique to measure vibration to assess the response of different latticed metastructures. 3D printed samples, each showing unique metastructure, are subjected to dynamic testing using a 3D Laser Doppler Vibrometer (LDV) from Polytec as shown in Fig. 6. 3D printed structures are made of PLA the detailed properties of the material and elemental composition of the hexagonal shape inserts made of mild steel are shown in supplementary Tables S3 and S4. These samples are attached to an Electrodynamic shaker system in cantilever mode. The laser head of LDV captures surface velocity by analyzing the phase shift between the incident and reflected laser beams. Data acquisition and signal processing are executed through the NI-DAQ system. The base excitation technique is utilized to determine transmissibility. In this method, the LDS shaker generates 3200 Fast Fourier Transform (FFT) lines, and the pseudo-random excitation with a frequency range of 0 to 10 KHz. Displacements of the top and bottom surfaces of the samples are recorded to calculate transmissibility after post-processing of the data. A retro-reflective tape is also attached to the surface where data has been recorded to reduce the scattering of the laser beam from the surface and three closely spaced points were taken on both the top and bottom surface for averaging of data to reduce error.
3.1 Performance measurement of lattice with resonators